# Assessing HRV and HR Dynamics with Wearables During Socially Anxious Situations: Insights from a Controlled Study in a Low-Middle-Income Country

### v1

@@ 4.5.1. Ablation Study
With male as the reference category, results in Table 5(a) show that females exhibit lower HRV, higher HR, and higher PAL than males. Moreover, to gain insights into the differences between participants with and without social anxiety within genders, we treated HRV/PAL as a dependent variable and phase (with baseline as reference), gender (with male as reference), and group (with non-SAD as reference) as fixed variables, and the remaining variables (participant id, program, and location) as random in the following model:
@@
The non-SAD group and the male gender were considered as the reference variables in the model. Consistent with our previous results, we observed that females consistently displayed lower HRV than their male counterparts shown in Table 5(b). Additionally, our results show that including gender as a fixed variable increased the number of distinguishing HRV parameters (including SDNN, RMSSD, S, SD1, SD2) between SAD and non-SAD groups. However, from hypothesis 2 testing, we found that only SDNN, RMSSD, and SD1 could significantly differentiate the SAD and non-SAD groups.
Table 5. Changes in HRV, HR, and Self-reported anxiety analysis with gender as a fixed variable. Where (i) Arrows, $\uparrow$ and $\downarrow$ represent increase and decrease, respectively, (ii) Significance levels ‘***’, ‘**’, ‘*’, ‘.’ correspond to 0.001, 0.01, 0.05, and 0.1, respectively
@@
5. Discussion
5.1. Negative correlation of HRV with self reported measures
@@
5.2. HR, HRV, and PAL across phases and their relationship between SAD and Non-SAD participants
We formulated three distinct hypotheses to address the following questions: (i) How does HRV and HR change across phases for all participants? (ii) What are the HRV and HR differences between participants with SAD and those without? (iii) How does self-reports (i.e., PAL) change across phases, regardless of group differences, and how does it differ between participants with SAD and those without?
@@ 5.2. HR, HRV, and PAL across phases and their relationship between SAD and Non-SAD participants
In Hypothesis 1, we examined the variations in HRV and HR during the anticipation, speech (activity), and reflection phases, irrespective of the participant groups (SAD vs. non-SAD). We observed notable changes in HRV and HR compared to the baseline in both phases (anticipation and speech), though the statistical significance of these changes exhibited variations between the phases (see Table 4(a)). In the anticipation phase, we found significantly lower SDNN, HTI, TINN, and ApEn values, but these differences did not maintain statistical significance in the speech phase. Conversely, in the speech phase, we found significant decreases in MeanNN, RMSSD, Prc20NN, pNN20, LnHF, SD1SD2, and an increase in HR, while these changes were not statistically significant during the anticipation phase. Only MedianNN, HF, and SD1 consistently exhibited significant reductions in both phases compared to the baseline. These findings indicate that while some HRV parameters showed consistent changes across both anticipation and activity phases, others exhibited phase-specific alterations.
@@
In the reflection phase, our observations revealed a mixture of outcomes, with notable variations in HRV features. Notably, we observed both significant increase and decrease in various HRV parameters. Specifically, MeanNN, MedianNN, Prc20NN, HF, HFn, and LnHF increased, signifying an augmentation in HRV during this phase. In contrast, HTI, TINN, and ApEn showed a decline. An interesting pattern emerged after excluding HTI, TINN, and ApEn from our analysis: HRV features that had exhibited lower values during the anticipation and activity phases demonstrated an increase in the reflection phase. This pattern suggests that participants may have felt more at ease during the reflection phase compared to the baseline and the preceding phases. This inference aligns with the observed increase in HRV levels and the decrease in HR during this phase, indicating a sense of physiological relaxation and reduced stress among the participants. However, our study methodology could not allow us to assess whether there was also a psychic change (cognition), e.g., self-criticism, being negatively evaluated, and cognitive distortions. Future research can incorporate questions or items on the current level of thinking patterns to gauge these changes. Moreover, a mixed-method study incorporating qualitative methods would be valuable in capturing these nuances of physiological or psychological changes.
In line with Hypothesis 2, we tried to understand the HRV patterns in both the SAD and non-SAD groups. Consistent with previous literature (alvares2013reduced, ; miranda2014anxiety, ; tolin2021psychophysiological, ; garcia2017autonomic, ; harrewijn2018heart, ; madison2021social, ), we indeed observed a consistent trend of lower HRV and higher HR in participants with SAD compared non-SAD (see Table 4(b)). However, it is worth noting that, contrary to the literature (gaebler2013heart, ; tolin2021psychophysiological, ; harrewijn2018heart, ), we found an insignificant increase in HF in SAD participants.
@@
Nevertheless, RMSSD and SD1 emerged as promising HRV parameters for distinguishing SAD participants from their non-SAD counterparts (see Table 4(b)), aligning with the findings of previous studies (alvares2013reduced, ; garcia2017autonomic, ; harrewijn2018heart, ; madison2021social, ). These parameters are potentially valuable physiological markers for identifying individuals with SAD. Although our results on HF contradict existing literature, it is essential to note that our findings align with the correlation results observed in Section 4.4, indicating that HF positively correlates with SPIN scores (SAD severity) during the speech phase of this study — however, this result regarding HF warrants further exploration to be fully validated.
@@
Furthermore, to understand the change in PAL with phases and between groups, we tested hypothesis 3. Similar to the existing literature (garcia2017autonomic, ; mauss2004there, ), we found that PAL significantly increased in anticipation and activity while there was a decrease (insignificant) in the reflection phase. These results imply that the anticipation and performance associated with tasks (speech) induced higher levels of anxiety compared to the baseline as shown in Figure 10. Furthermore, the analysis revealed a significant difference in PAL between the SAD and non-SAD groups. This finding suggests that individuals with SAD perceive and report higher levels of anxiety than those without SAD. Figure 11 shows the changes in the SDNN, RMSSD, SD1, HR, and the PAL across the baseline, anticipation, speech, and reflection.
5.3. Effect of Gender on HR and HRV
@@
In addition to the correlation and hypothesis testing, we explored how HRV and HR vary between males and females. The Figure 12 shows the changes in the HR of males and females across the baseline, anticipation, speech, and reflection. In alignment with existing research (geovanini2020age, ), our findings indicated that, as a general trend, females demonstrated a higher HR and lower HRV levels in comparison to males. Furthermore, we discovered that both males and females with SAD displayed lower HRV than their non-SAD counterparts within their respective gender groups (alvares2013reduced, ).
@@
Our correlation analysis revealed that gender influences HRV and PAL. Based on this, we included gender as a fixed variable in our analysis to understand its effect on HRV while treating groups as a fixed variable. We found that by including gender as a fixed variable, the HRV parameters S and SD2 can also serve as physiological markers for SAD. These gender-specific insights deepen our understanding of HRV patterns and differences related to SAD among males and females. This finding is novel and has not been reported in existing literature, warranting further exploration.
5.4. Physiological markers of SAD
@@
Based on our findings and the existing literature, we advocate for a comprehensive exploration of HRV features beyond HF and RMSSD. In particular, we propose that SDNN, RMSSD, and SD1 as the potential physiological markers of SAD. Our study is notable for being the first to identify SD1 (nonlinear HRV), representing short-term RR interval fluctuations, as a potential physiological marker for distinguishing participants with SAD from those without SAD. Additionally, SD1 played a pivotal role in unveiling the dynamic changes in HRV during the different phases of our study. Specifically, SD1 indicates that HRV decreases during the anticipation and activity phases while increasing during the reflection phase compared to the baseline.
Moreover, when considering gender as a factor, TINN, S, and SD2 also emerge as potential physiological markers of SAD (see Table 5(a)). However, further exploration is needed to validate these physiological markers, as we are the first to propose them as indicators of anxiety disorders with gender as a factor. Examining these features can provide deeper insights into the physiological responses of individuals with SAD and contribute to a more holistic understanding of HRV alterations in gender specific SAD population.
Figure 11. Differences in physiological measures (SDNN, RMSSD, SD1, HR) and perceived anxiety level (PAL) across the baseline (B), anticipation (ATP), speech activity (AP), and reflection (RP) phases between SAD and non-SAD participants. [Best viewed in color]
@@
Figure 12. Differences in HR between different groups. [Best viewed in color]
Figure 13. Mean PAL of male and female participants at different phases (Baseline, Anticipation, Speech, and Reflection) of the study. [Best viewed in color]
5.5. Implications for mental health researchers
@@
Ideally, any physiological change in the participants should correspond to their PAL. For instance, in continuation with changes observed in above PAL, the decrease in HRV should be more prominent during the anticipation phase, followed by the speech phase, and conversely, an opposing trend should be observed during the reflection phase. To this end, we observed changes in HF in sync with the PAL. HF, an HRV parameter extensively explored in literature, exhibited a significant reduction ($p<0.01$) in anticipation compared to the baseline (see Table 4(a)), suggesting that the participants were in a relaxed state in the baseline. Furthermore, there was a reduced HF during speech, although the significance level dropped to 0.1, indicating a decreased difference between the HF of speech and baseline. Notably, during reflection, HF increased compared to baseline. The changes observed in HF align with the changes in PAL, confirming that HF is the most reliable predictor of “anxious state”. However, it is crucial to note that this evaluation of HF is irrespective of SAD/non-SAD groups.
Moreover, the conclusion about HF could be debated as the pattern of changes in HF and the PAL did not exhibit consistency with other HRV measures. For instance, with MeanNN, MedianNN, Prc20NN, SD1, SD1SD2, and HR, the notable differences were found to be higher during the speech activity rather than during the anticipation of the activity (see Table 4(a)). Yet, it is crucial to note that the HRV measures (MeanNN, MedianNN, Prc20NN, and HR) highly correlated with PAL in activity, while HF exhibited a higher correlation with PAL during the anticipation phase.
@@
The analysis on the gender reveals that females consistently demonstrated significantly lower heart rate variability (MeanNN, MedianNN, SDNN, RMSSD, HTI, S, SD1, SD2) and higher heart rate levels throughout the study than males (see Table 5(a)). Nonetheless, it is crucial to highlight that in comparison to the male group, the female group exhibited lower PAL during the baseline and anticipation phases but reported higher PAL during anticipation and reflection (see Figure 13). One possible explanation for this could be linked to higher self-esteem among females (students in our study), as reported in various studies (tamanaifar2023relationship, ). Additionally, it is important to note that anxious females exhibited decreased HRV and elevated heart rate compared to non-anxious females.
5.5.1. Wearables and data analysis
@@
  • Firstly, our analysis indicates that RMSSD, SDNN, and SD1 are reliable physiological markers for capturing social anxiety. However, our findings are based on data collected with clinical-grade ECG sensors. It is important to note that PPG is easily sensed by smartwatches and serves as a proxy for ECG (jaiswal2023comparative, ; kinnunen2020feasible, ; theurl2023smartwatch, ). Therefore, PPG should be further explored to validate our findings and extend the use of these physiological markers for SAD detection in real-world settings.
@@
5.6. Limitations
Our study has several important limitations: (i) Our study primarily involved non-clinical participants (institute students). The SAD and non-SAD categories were grouped based on the SPIN score. Though SPIN is a validated self-reported questionnaire for finding social anxiety in participants, it may not be able to capture the clinically significant SAD. (ii) Data loss: The study was conducted with ninety-nine participants (SAD - 71 and non-SAD - 28), but forty-eight participants (SAD-40 and non-SAD-8) data were dropped due to noise and artifacts. However, ECG data loss is common in cardiovascular research due to noise and artifacts (held2021heart, ). The data loss may or may not have affected the results. Future studies should consider implementing thorough skin preparation and enhancing electrode-skin contact to mitigate such issues. These limitations emphasize the need for future research to address these issues to comprehensively understand SAD and its physiological correlates.
5.7. Future work

### v2

@@ 4.5.1 Ablation Study
With male as the reference category, results in Table 5(a) show that females exhibit lower HRV, higher HR, and higher PAL than males irrespective of phases. Moreover, to gain insights into the differences between participants with and without social anxiety within genders, we treated HRV/PAL as a dependent variable and phase (with baseline as reference), gender (with male as reference), and group (with non-SAD as reference) as fixed variables, and the remaining variables (participant id, program, and location) as random in the following model:
@@
Consistent with our previous results, we observed that females consistently displayed lower HRV than their male counterparts shown in Table 5(b). Additionally, our results show that including gender as a fixed variable increased the number of distinguishing HRV parameters (including SDNN, RMSSD, TINN, S, SD1, and SD2) between SAD and non-SAD groups. However, from hypothesis 2 testing, we found that only SDNN, RMSSD, S and SD1 could significantly differentiate the SAD and non-SAD groups.
Table 5: Changes in HRV, HR, and Self-reported anxiety analysis with gender as a fixed variable. Where (i) Arrows, $\uparrow$ and $\downarrow$ represent increase and decrease, respectively, (ii) Significance levels ‘***’, ‘**’, ‘*’, ‘.’ correspond to 0.001, 0.01, 0.05, and 0.1, respectively
@@
5 Discussion
5.1 Negative correlation of HRV with self-reported measures
@@
5.2 HR, HRV, and PAL across phases and their relationship between SAD and Non-SAD participants
To answer (i) How does HRV and HR change across phases for all participants? (ii) What are the HRV and HR differences between participants with SAD and those without? (iii) How does self-reports (i.e., PAL) change across phases, regardless of group differences, and how does it differ between participants with SAD and those without?, we hypothesized: Hypothesis 1 (H1): There will be a significant increase in HR and a decrease in HRV during the Anticipation and Speech activity phases, respectively, compared to the Baseline for all participants. Additionally, HR and HRV will exhibit significant differences during the Reflection phase compared to the Baseline phase for all participants. Hypothesis 2 H2): There will be significant differences in HR and HRV between the SAD and non-SAD participants during the anxious activity. Hypothesis 3 (H3): There will be a significant increase in perceived anxiety levels during the Anticipation and Speech activity phases compared to the Baseline for all participants. Furthermore, perceived anxiety levels during the Reflection phase will significantly differ from those during the Baseline.
@@ 5.2 HR, HRV, and PAL across phases and their relationship between SAD and Non-SAD participants
In H1, we examined the variations in HRV and HR during the anticipation, speech (activity), and reflection phases, irrespective of the participant groups (SAD vs. non-SAD). We observed notable changes in HRV and HR compared to the baseline in both phases (anticipation and speech). However, the statistical significance of these changes exhibited variations between the phases (see Table 4(a)). In the anticipation phase, we found significantly lower SDNN, HTI, TINN, and ApEn and significantly higher HF, HFn, and LnHF values. However, the statistical significance of these differences changed in the speech phase. Similarly, in the speech phase, we found significant decreases in MeanNN, MedianNN, RMSSD, Prc20NN, pNN20, LnHF, S, SD1, SD1SD2, and an increase in HR. Only MedianNN, RMSSD, TINN, S, and SD1 consistently exhibited significant reductions in both anticipation and speech phases compared to the baseline. These findings indicate that while some HRV parameters showed consistent changes across both anticipation and activity phases, others exhibited phase-specific alterations.
@@
In the reflection phase, our observations revealed a mixture of outcomes with notable variations in HRV features. Notably, we observed both significant increases and decreases in various HRV parameters. Specifically, MeanNN, MedianNN, Prc20NN, HF, HFn, and LnHF increased, signifying an augmentation in HRV during this phase. In contrast, HTI, TINN, and ApEn showed a decline. An interesting pattern emerged after excluding HTI, TINN, ApEn, and frequency domain HRV from our analysis: HRV features that had exhibited lower values during the anticipation and activity phases demonstrated an increase in the reflection phase. This pattern suggests that participants may have felt more at ease during the reflection phase compared to the baseline and the preceding phases. This inference aligns with the observed increase in HRV levels and the decrease in HR during this phase, indicating a sense of physiological relaxation and reduced stress among the participants. However, our study methodology could not allow us to assess whether there was also a psychic change (cognition), e.g., self-criticism, being negatively evaluated, and cognitive distortions. Future research can incorporate questions or items on the current level of thinking patterns to gauge these changes. Moreover, a mixed-method study incorporating qualitative methods would be valuable in capturing these nuances of physiological or psychological changes.
@@
In line with Hypothesis 2, we tried to understand the HRV patterns in both the SAD and non-SAD groups. Consistent with previous literature [alvares2013reduced, miranda2014anxiety, tolin2021psychophysiological, garcia2017autonomic, harrewijn2018heart, madison2021social, gaebler2013heart, tolin2021psychophysiological, harrewijn2018heart], we indeed observed a consistent trend of lower HRV and higher HR in participants with SAD compared non-SAD (see Table 4(b)). Nevertheless, RMSSD, S, and SD1 emerged as promising HRV parameters for distinguishing SAD participants from their non-SAD counterparts (see Table 4(b)), aligning with the findings of previous studies [alvares2013reduced, garcia2017autonomic, harrewijn2018heart, madison2021social]. These parameters are potentially valuable physiological markers for identifying individuals with SAD. Moreover, our findings on HF are consistent with existing literature [alvares2013reduced, pittig2013heart, gaebler2013heart, tolin2021psychophysiological, harrewijn2018heart]; however, the results in our study were not statistically significant. The correlation analysis revealed that HF was negatively associated with SPIN and PAL during the baseline, anticipation, and reflection phases but showed a positive correlation during the activity phase. This shift suggests that the role of HF may vary across different phases, indicating the need for further investigation.
@@
Furthermore, to understand the change in PAL with phases and between groups, we tested hypothesis 3. Similar to the existing literature [garcia2017autonomic, mauss2004there], we found that PAL significantly increased in anticipation and activity while there was a decrease (insignificant) in the reflection phase. These results imply that the anticipation and performance associated with tasks (speech) induced higher levels of anxiety compared to the baseline as shown in Figure 10. Furthermore, the analysis revealed a significant difference in PAL between the SAD and non-SAD groups. This finding suggests that individuals with SAD perceive and report higher levels of anxiety than those without SAD. Figure 11 shows the changes in the SDNN, RMSSD, SD1, HR, and the PAL across the baseline, anticipation, speech, and reflection.
5.3 Effect of Gender on HR and HRV
@@
In addition to the correlation and hypothesis testing, we explored how HRV and HR vary between males and females. The Figure 12 shows the changes in the HR of males and females across the baseline, anticipation, speech, and reflection. In alignment with existing research [geovanini2020age], our findings indicated that, as a general trend, females demonstrated a higher HR and lower HRV levels in comparison to males. Furthermore, we discovered that both males and females with SAD displayed lower HRV than their non-SAD counterparts within their respective gender groups [alvares2013reduced].
@@
Our correlation analysis revealed that gender influences HRV and PAL. Based on this, we included gender as a fixed variable in our analysis to understand its effect on HRV while treating groups as a fixed variable. We found that by including gender as a fixed variable, the HRV parameters TINN and SD2 can also serve as physiological markers for SAD. However, the significance level was only 0.1. These gender-specific insights deepen our understanding of HRV patterns and differences related to SAD among males and females. This finding is novel and, to our knowledge, has not been reported in existing literature, warranting further exploration. Moreover, our analysis of gender requires further exploration due to the low number of female participants (33%) compared to male participants.
5.4 Physiological markers of SAD
@@
Based on our findings and the existing literature, we advocate for a comprehensive exploration of HRV features beyond HF and RMSSD. In particular, we propose that RMSSD, S, and SD1 as the potential physiological markers of SAD. Our study is notable for being the first to identify SD1 (nonlinear HRV), representing short-term RR interval fluctuations, as a potential physiological marker for distinguishing participants with SAD from those without SAD. Additionally, SD1 played a pivotal role in unveiling the dynamic changes in HRV during the different phases of our study. Specifically, they indicate that HRV decreases during the anticipation and activity phases while increasing during the reflection phase compared to the baseline.
Moreover, when considering gender as a factor, TINN and SD2 also emerge as potential physiological markers of SAD (see Table 5(a)). However, further exploration is needed to validate these physiological markers, as we are the first to propose them as indicators of anxiety disorders with gender as a factor. Examining these features can provide deeper insights into the physiological responses of individuals with SAD and contribute to a more holistic understanding of HRV alterations in gender-specific SAD populations.
Figure 11: Differences in physiological measures (SDNN, RMSSD, SD1, HR) and perceived anxiety level (PAL) across the baseline (B), anticipation (ATP), speech activity (AP), and reflection (RP) phases between SAD and non-SAD participants. [Best viewed in color]
@@
Figure 12: Differences in HR between different groups. [Best viewed in color]
5.5 Implications for mental health researchers
@@
Figure 13: Mean PAL of male and female participants at different phases (Baseline, Anticipation, Speech, and Reflection) of the study. [Best viewed in color]
Ideally, any physiological change in the participants should correspond to their PAL. For instance, in continuation with changes observed in the above PAL, the decrease in HRV should be more prominent during the anticipation phase, followed by the speech phase, and conversely, an opposing trend should be observed during the reflection phase. To this end, we observed changes in MeanNN, MedianNN, Prc20NN, SD1, and SD1SD2 in sync with the PAL. However, it is crucial to note that these findings are irrespective of SAD/non-SAD groups.
@@
The analysis on the gender reveals that females consistently demonstrated significantly lower heart rate variability (MeanNN, MedianNN, SDNN, RMSSD, HTI, S, SD1, SD2) and higher heart rate levels throughout the study than males (see Table 5(a)). Nonetheless, it is crucial to highlight that in comparison to the male group, the female group exhibited lower PAL during the baseline and anticipation phases but reported higher PAL during anticipation and reflection (see Figure 13). One possible explanation for this could be linked to higher self-esteem among females (students in our study), as reported in various studies [tamanaifar2023relationship]. Additionally, it is important to note that anxious females exhibited decreased HRV and elevated heart rate compared to non-anxious females.
5.5.1 Wearables and data analysis
@@
  • Firstly, our analysis indicates that RMSSD, SDNN, and SD1 are reliable physiological markers for capturing social anxiety. However, our findings are based on data collected with clinical-grade ECG sensors. It is important to note that PPG is easily sensed by smartwatches and serves as a proxy for ECG [jaiswal2023comparative, kinnunen2020feasible, theurl2023smartwatch]. Therefore, PPG should be further explored to validate our findings and extend the use of these physiological markers for SAD detection in real-world settings.
@@
5.6 Contextual relevance of our findings
Most existing studies on physiological markers for anxiety disorders have been conducted in the Global North [alvares2013reduced, gaebler2013heart, tolin2021psychophysiological, miranda2014anxiety, tamura2013salivary, garcia2017autonomic, madison2021social, held2021heart, harrewijn2018heart, bailey2019moderating], where healthcare infrastructure is more advanced and mental health services are more accessible. However, the experience and expression of mental disorders are strongly influenced by cultural beliefs, social norms, and systemic stigma, all of which can differ significantly across regions [pendse2019mental]. Factors such as gender roles, social networks, and geopolitics further shape how individuals perceive and respond to mental health challenges [wani2024unrest].
In contrast, the Global South—particularly countries like India—faces unique challenges, including widespread stigma, limited access to mental health services, and a lack of culturally grounded research [pendse2019mental]. Our study addresses this gap by examining physiological and self-reported markers of SAD within an Indian context. To our knowledge, this is the first study in India to investigate HR and HRV variations during anxiety-inducing activities using wearable sensors. India, the world’s most populous country, has a highly diverse population spanning multiple cultures, languages, and socioeconomic backgrounds [bose2021integrating]. The participants in our study were students studying at IISER Bhopal, a premier academic institution that draws students from various regions across India. This diversity allowed us to capture various responses across urban and rural settings and genders.
By analyzing both objective physiological data and subjective self-reports from participants with and without SAD, our study provides insights into how anxiety manifests when participants anticipate, perform, and reflect on an anxiety-provoking activity. These findings are essential for designing mental health interventions that are not only effective but are also context-aware. The publicly available dataset generated through this work contributes a rare resource for future research focused on LMICs. It offers a foundation for building localized and inclusive mental health technologies, aligning with the COMPASS vision of computing that is situated, socially responsive, and driven by the realities of place-based challenges.
6 Limitations and Future work
Our study has several important limitations: (i) Our study primarily involved non-clinical participants (institute students). The SAD and non-SAD categories were grouped based on the SPIN score. Though SPIN is a validated self-reported questionnaire for finding social anxiety in participants, it may not be able to capture the clinically significant SAD. (ii) Data loss: The study was conducted with ninety-nine participants (SAD - 71 and non-SAD - 28), but forty-eight participants (SAD-40 and non-SAD-8) data were dropped due to noise and artifacts. However, ECG data loss is common in cardiovascular research due to noise and artifacts [held2021heart]. The data loss may or may not have affected the results. Future studies should consider implementing thorough skin preparation and enhancing electrode-skin contact to mitigate such issues. These limitations emphasize the need for future research to address these issues to comprehensively understand SAD and its physiological correlates.